. An old fence used to enclose a circular lot is to be recycled to fence a square lot. if the radius of a circular lot is 5 meters, how long should the edge of the square lot be?
1
Expert's answer
2012-08-10T11:27:49-0400
Perimeter of a fence is

P = 2*π*5 ≈ 31.4160 m.

So, the edge of the square lot should be

L = P/4 = 7.8540 m.